Inner radius of sperical sheel is 3.7cm and outer radius is 4.5 cm and volume charge density is 6.1 x 10^-4 c/m^3 what is the total charge ?
1
Expert's answer
2020-03-27T10:55:08-0400


So, we have


"q=\Delta V\cdot\rho=(\frac{4}{3}\pi r_2^3-\frac{4}{3}\pi r_1^3)\cdot \rho=\frac{4}{3}\pi (r_2^3-r_1^3)\cdot \rho="


"=\frac{4}{3}\cdot 3.14\cdot (0.045^3-0.037^3)\cdot 6.1\cdot10^{-4}=1.67\cdot10^{-6}C=1.67" "\mu C"